OpenStudy (18jonea):

Find the area of the entire slice of pizza. Since the slice of pizza is part of a whole pie, let’s think of the pizza as a sector of a circle. Calculate the area. Be sure to show all of you work.

OpenStudy (18jonea):



OpenStudy (18jonea):

@welshfella @amistre64

OpenStudy (18jonea):

@dan815

OpenStudy (welshfella):

the radius of the circle is 6 so area of the whole pizza is pi * 6^2 = ? now you need to find out how many pieces of pizza are in the whole pizza this equals 360/45 = x so divide the area of the whole pizza by x and you'll get it.

OpenStudy (18jonea):

113.04 is area of whole pizza

OpenStudy (18jonea):

8 pieces

OpenStudy (18jonea):

so the answer would be 14.13

OpenStudy (welshfella):

right so area of 1 pizza clice = 113.04 / 8

OpenStudy (18jonea):

14.13

OpenStudy (welshfella):

right

OpenStudy (18jonea):

then I have to do this for my next part of the questionFind the area of the pizza without the crust. We can use our knowledge of triangles for this step. Given that the height of the triangle is 5.5 in. and the base of the triangle is 5 in., calculate the area. Show all of you work

OpenStudy (welshfella):

what did you use for pi , 3.14?

OpenStudy (18jonea):

i got the answer too Area of a triangle = ½ base x height Area of triangle = ½ 5 x 5.5 Area of triangle = 13.75 in squared

OpenStudy (welshfella):

looks good

OpenStudy (18jonea):

thn can you help me with the are of the crust? would i just subtract 13.75 from 14.13 to get the answer?

OpenStudy (18jonea):

yes i used 3.14 for pi

OpenStudy (welshfella):

yes

OpenStudy (18jonea):

so the are is only .38?

OpenStudy (welshfella):

right

OpenStudy (18jonea):

that seems really small though

OpenStudy (welshfella):

i got 113.097 for the area of whole pizza because i used pi on my calculator but if you usually use 3.14 in class thats ok.

OpenStudy (welshfella):

thin crust! well you have your calculations right

OpenStudy (18jonea):

ok thank you

OpenStudy (welshfella):

yw